On January 1, 2013, the Mason Manufacturing Company began construction of a building to be used as its office headquarters. Ti want to know the interest capitalized and interest expense in 2019 and the cost of the building.
Answer #1

If it is helpful, please rate the answer and if any doubt arises let me know

2018 2019
Interest Capitalized $               474,000 $   582,819
Interest Expense $            1,016,000 $   907,181
Cost of Building $            7,835,819
Workings
Expenditure for 2018
Jan 1,2018 $            13,50,000 X 12/12 = $        13,50,000
March 1, 2018 $            10,80,000 X 10/12 = $          9,00,000
June 30, 2018 $            12,80,000 X 6/12 = $          6,40,000
October 1, 2018 $            10,80,000 X 3/12 = $          2,70,000
$            47,90,000 $        31,60,000
Interest Capitalized in 2018
$            31,60,000 X 15.00% = $          4,74,000
Expenditure for 2019
Jan 1, 2019 ($4790000+$474,000) $            52,64,000 X 9/9 = $        52,64,000
Jan 31, 2019 $               3,42,000 X 8/9 = $          3,04,000
April 30, 2019 $               6,75,000 X 5/9 = $          3,75,000
August 31. 2019 $               9,72,000 X 1/9 = $          1,08,000
$            72,53,000 $        60,51,000
Interest Capitalized in 2019
$                      60,51,000
Less: $                      38,00,000 X 15.00% X 9/12 = $          4,27,500
$                      22,51,000 X 9.20% X 9/12 = $          1,55,319
Interest Capitalized in 2019 = $          5,82,819
Weighted Average rate of all debt:-
$            40,00,000 X 8% = $          3,20,000
$            60,00,000 X 10% = $          6,00,000
$         1,00,00,000 $          9,20,000
Weighted Average rate of all debt = 9.20%
($ 920,000/ $ 100,00,000)
Interest Expense for 2018:
Total Interest Incurred ($38,00,000*15%)+($40,00,000*8%)+($60,00,000*10%) = $        14,90,000
Less : Interest Capitalized = $          4,74,000
2018 Expense = $        10,16,000
Interest Expense for 2019:
Total Interest Incurred = $        14,90,000
Less : Interest Capitalized = $          5,82,819
2019 Expense = $          9,07,181
Cost of Building
Expenditure for 2018 = $        47,90,000
Interest Capitalized in 2018 = $          4,74,000
Expenditure for 2019 = $        19,89,000
Interest Capitalized in 2019 = $          5,82,819
Cost of Building = $        78,35,819